# From Word to World: Evaluate and Mitigate Culture Bias via Word Association Test

### v1

@@ 5.2 Experimental Setup
5.3 Overall Results
LLMs Exhibit Cultural Biases in the WAT Task
The overall comparison results are presented in Table 1, where the “Baseline” denotes the default cultural preferences exhibited by the LLMs. Both models consistently reveal a cultural preference pattern: USA > UK > OC > CN. [^6] To demonstrate that performance differences are due to cultural factors rather than the language used in prompts, we conducted an ablation study. Details are provided in Appendix C. This suggests that, in terms of word-level cognitive abilities, LLMs exhibit a stronger inclination toward American culture, aligning with findings from prior studies Cao et al. (2023b); Myung et al. (2024); Zhou et al. (2025). Overall, Llama demonstrates stronger human-like word-level cognitive abilities compared to Qwen, primarily due to its significantly better performance in English-speaking cultural contexts. In contrast, the performance gap between the two models is relatively small in CN. Interestingly, when evaluating PWR@3, Qwen outperforms Llama across all cultural contexts, suggesting that Qwen excels at predicting highly associated words. However, as K increases, Llama surpasses Qwen, indicating that Llama possesses a more comprehensive associative cognition across a broader range of related words.
CultureSteer Enhance LLM Cultural Awareness
At the methodological level, CultureSteer consistently outperforms both prompt-based methods and the base models (Llama and Qwen) across all cultural settings. This demonstrates that the model does not rely on explicit textual configurations. Instead, it leverages linear transformations to implicitly align with the semantic space of specific cultural preferences, effectively adjusting the default semantic space to enhance its cultural awareness. Interestingly, for the Llama model, the CSP method consistently outperforms the CCT method across all cultural contexts. In contrast, for the Qwen model, the CSP method performs better only in the Chinese cultural context, while the CCT method excels in the other three, particularly in English-speaking settings. This suggests that Llama is more effective with culture-specific prompts tailored to a single context, whereas Qwen demonstrates stronger cross-cultural reasoning, especially for the English-speaking environments. This demonstrates that explicitly simulating cultural contexts through prompt design highlights the sensitivity of LLM performance to specific prompt configurations. Moreover, it further validates the effectiveness of CultureSteer’s implicit learning paradigm in enhancing cultural awareness.
5.4 Fine-Grained Results
6 Analysis
6.1 Culture-Specific Semantic Learning Matters
@@ A.3 Data Overview
Chap Cue Asso (avg.) Users
@@
Appendix B Experiment Settings
B.1 Template Using
@@
EN,CN-t Base When "{cue_word}" is mentioned, people often think of the following words:
CSP You are a person with {culture} cultural background. You will be performing a word association task. Please directly answer the association word. When "{cue_word}" is mentioned, people often think of the following words:
CCT You are a person with {culture} cultural background. You will be performing a word association task. Please directly answer the association word. Before you respond, think about how {culture} culture is different from {cultures} cultures. When "{cue_word}" is mentioned, people often think of the following words
CN Base {CJK*}UTF8gbsn当提起"{cue_word}",人们往往会想到的词是:
CSP {CJK*}UTF8gbsn你是一个中国文化背景的人。 {CJK*}UTF8gbsn你将进行词联想任务,请直接说出你联想到的词。 {CJK*}UTF8gbsn当提起"{cue_word}",人们往往会想到的词是:
CCT {CJK*}UTF8gbsn你是一个中国文化背景的人。 {CJK*}UTF8gbsn你将进行词联想任务, 请直接说出你联想到的词。 {CJK*}UTF8gbsn在回答之前, 请你注意中国文化与美国、英国、大洋洲文化的不同。 {CJK*}UTF8gbsn当提起"{cue_word}",人们往往会想到的词是:
@@
B.2 Hyperparaments
The detailed setting of hyperparaments are shown in Table  7
@@ B.2 Hyperparaments
Appendix C Language Ablation Study
@@
Appendix D Granular Performance
The granular performance is shown in Figure 5. When K=3, 5, 10, the CultureSteer results consistently exhibit comparably weaker performance in Global Knowledge categories, while cultural disparities in Llama remain greater than those in Qwen.
(a) Performance in PWR@3
(b) Performance in PWR@5
(c) Performance in PWR@10
Figure 5: Graunalr performance in PWR@3, 5 and 10

### v2

@@ 5.3 Overall Results
5.4 Fine-Grained Results
6 Analysis
6.1 Culture-Specific Semantic Learning Matters
@@ 7 Conclusion
Acknowledgments
This work is supported by Shenzhen Science and Technology Program (Shenzhen Key Laboratory, Grant No. ZDSYS20230626091302006), Program for Guangdong Introducing Innovative and Entrepreneurial Teams (Grant No. 2023ZT10X044), Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing, Shenzhen Science and Technology Research Fund (Fundamental Research Key Project, Grant No.JCYJ20220818103001002), the International Science and Technology Cooperation Center, Ministry of Science and Technology of China (under grant 2024YFE0203000), the Internal Project of Shenzhen Research Institute of Big Data.
@@ A.3 Data Overview
Chap Cue
Asso
Users
@@
Appendix B Metric Discussing
We provide a detailed comparison with Discounted Cumulative Gain at K (DCG@K), a widely-used position-weighted evaluation metric in information retrieval.
DCG@K Formulation
DCG@K applies a logarithmic discount to the relevance scores based on their ranking positions, with the intuition that items ranked lower should contribute less to the overall score. For our Word Association Task (WAT), we adapt DCG@K as follows. Given the ground truth associative words $A_{w}^{c}=\{a_{1},a_{2},\ldots,a_{N}\}$ ranked by their associative strength, and the LLM’s predicted ranking $\hat{A}_{w}$, DCG@K is defined as:
$\text{DCG@K}=\sum_{i=1}^{K}\frac{\text{rel}(i)}{\log_{2}(i+1)}$ (8)
where $\text{rel}(i)$ represents the relevance score of the item at position $i$ in the predicted ranking. In our adaptation, we define:
$\text{rel}(i)=\begin{cases}\frac{N-\text{rank}(\hat{a}_{i})+1}{N}&\text{if }\hat{a}_{i}\in A_{w}^{c}\\ 0&\text{otherwise}\end{cases}$ (9)
where $\text{rank}(\hat{a}_{i})$ denotes the ground truth ranking position of the predicted word $\hat{a}_{i}$ in $A_{w}^{c}$. This formulation assigns higher relevance scores to words that appear earlier in the ground truth ranking.
Experimental Results with DCG@K
We re-evaluate all our main experiments using DCG@K and present the results in Table 6. The experimental setup and datasets remain identical to those reported in the main paper.
The results obtained using NDCG@K are consistent with those derived from our proposed PWR@K metric. Both metrics demonstrate similar trends across different models, cultural contexts, and experimental conditions. This consistency validates that our key findings and conclusions are robust to the choice of position-weighted evaluation metric.
Llama Qwen
PWR@K K=3 K=5 K=10 K=20 K=3 K=5 K=10 K=20
[…]
CultureLLM 0.78 0.93 1.08 1.12 0.29 0.35 0.44 0.54
SimLLMCultureDist 0.98 1.27 1.63 1.87 0.82 1.00 1.23 1.51
CultureMerge 0.92 1.20 1.57 1.79 - - - -
CultureSPA 0.87 1.15 1.53 1.76 - - - -
CSP 1.04 1.31 1.65 1.90 1.02 1.22 1.50 1.77
CCT 0.83 1.06 1.33 1.49 1.04 1.21 1.46 1.72
USA CultureSteer 1.49 1.87 2.36 2.66 1.38 1.71 2.13 2.40
[…]
CultureLLM 0.70 0.81 0.85 0.62 0.25 0.30 0.37 0.45
SimLLMCultureDist 1.03 1.28 1.43 0.77 0.69 0.82 1.00 1.20
CultureMerge 0.97 1.22 1.38 0.74 - - - -
CultureSPA 0.95 1.20 1.37 0.68 - - - -
CSP 1.00 1.24 1.41 0.81 0.79 0.95 1.16 1.35
CCT 0.85 1.04 1.17 0.68 0.81 0.95 1.15 1.33
UK CultureSteer 1.38 1.71 1.94 1.16 1.03 1.27 1.55 1.77
[…]
CultureLLM 0.62 0.72 0.83 0.86 0.28 0.33 0.40 0.49
SimLLMCultureDist 0.77 0.98 1.22 1.39 0.70 0.83 1.01 1.20
CultureMerge 0.74 0.95 1.20 1.37 - - - -
CultureSPA 0.68 0.90 1.17 1.32 - - - -
CSP 0.81 1.00 1.22 1.39 0.82 0.98 1.19 1.37
CCT 0.68 0.85 1.04 1.16 0.79 0.93 1.10 1.28
OC CultureSteer 1.16 1.44 1.77 1.99 1.04 1.28 1.56 1.79
[…]
CultureLLM 0.06 0.06 0.07 0.07 0.29 0.38 0.52 0.65
SimLLMCultureDist 0.58 0.76 0.96 1.11 0.77 0.98 1.26 1.58
CultureMerge 0.59 0.76 0.95 1.08 - - - -
CultureSPA 0.62 0.79 1.00 1.14 - - - -
CSP 0.80 1.03 1.30 1.47 1.13 1.45 1.94 2.35
CCT 0.81 1.01 1.24 1.39 0.94 1.24 1.68 2.06
CN CultureSteer 1.06 1.35 1.69 1.87 1.45 1.88 2.54 3.06
[…]
CultureLLM 0.71 0.60 0.70 0.73 0.33 0.32 0.40 0.50
SimLLMCultureDist 0.99 0.97 1.22 1.41 0.85 0.86 1.06 1.29
CultureMerge 0.95 0.93 1.19 1.37 - - - -
CultureSPA 0.90 0.91 1.17 1.36 - - - -
CSP 1.02 1.03 1.29 1.50 1.00 1.07 1.33 1.61
CCT 0.84 0.89 1.11 1.27 0.98 1.01 1.24 1.50
Average CultureSteer 1.45 1.44 1.80 2.07 1.31 1.43 1.78 2.13
Table 6: Re-evaluated experimental results using DCG@K show consistent performance trends with our proposed PWR@K metric across models and conditions, confirming the validity and robustness of PWR@K for assessing ranking effectiveness.
Appendix C Experiment Settings
C.1 Template Using
@@
EN,CN-t Base When ”{cue_word}” is mentioned, people often think of the following words:
CSP
You are a person with {culture} cultural background. You will be performing a word association task.
Please directly answer the association word.
When ”{cue_word}” is mentioned, people often think of the following words:
CCT
You are a person with {culture} cultural background. You will be performing a word association task.
Please directly answer the association word.
Before you respond, think about how {culture} culture is different from {cultures} cultures.
When ”{cue_word}” is mentioned, people often think of the following words
CN Base 当提起”{cue_word}”,人们往往会想到的词是:
CSP
你是一个中国文化背景的人。
你将进行词联想任务,请直接说出你联想到的词。
当提起”{cue_word}”,人们往往会想到的词是:
CCT
你是一个中国文化背景的人。 你将进行词联想任务, 请直接说出你联想到的词。
在回答之前, 请你注意中国文化与美国、英国、大洋洲文化的不同。
当提起”{cue_word}”,人们往往会想到的词是:
@@
C.2 Hyperparaments
The detailed setting of hyperparaments are shown in Table 8.
@@ C.2 Hyperparaments
Appendix D Language Ablation Study
@@
Appendix E Granular Performance
The granular performance is shown in Figure 7(a). When K=3, 5, 10, the CultureSteer results consistently exhibit comparably weaker performance in Global Knowledge categories, while cultural disparities in Llama remain greater than those in Qwen.
(a) Graunalr performance in PWR@3, 5 and 10.